Gabriele Enterprises has bonds on the market making annual payments, with 16 years to maturity, a par value of $1,000, and selling for $880. At this price, the bonds yield 11 percent. What must the coupon rate be on the bonds?
Answer #1

Current price=Annual coupon*Present value of annuity factor(11%,16)+$1000*Present value of discounting factor(11%,16)

880=Annual coupon*7.37916178+$1000*0.188292204

Annual coupon=(880-188.292204)/7.37916178

=93.74(Approx).

Coupon rate=Annual coupon/Par value

=(93.74/1000)

=9.37%(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=Annual coupon[1-(1.11)^-16]/0.11

=Annual coupon*7.37916178

2.Present value of discounting factor=1000/1.11^16

=1000*0.188292204